A baseball team has home games on Wednsday and Sunday. The two games together earn $5041.00 for the team. Wednesday's game generates $1409.00 less than Sunday's game. How much money was taken in at each game.
Answer by addingup(3677)

x + x-1409 = 5041
2x = 6450
x = 3225 this was Sunday
Wednesday: 3225-1409 = 1816 this was Wednesday.
--------------------
Check:
3225 + 1816 = 5041 correct